# Inattention to States and Characteristics

### v1

@@ Conclusion
The exchangeability of the integral and the derivative follows from [3, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$. The result follows by construction of $Y$. ∎
@@
Proof of corollary˜11.1.
Suppose $\gamma,\nu$ are two marginals, and without loss $V(\gamma)\geq V(\nu)$. We can upper bound $V(\gamma)$ by using the “wrong” Schrödinger potentials in the dual objective. In particular, let $\gamma=\nu_{\varepsilon}$
$|V(\nu_{\varepsilon})-V(\nu)|\leq\bigg{|}\int a_{\nu}\>d(\nu_{\varepsilon}-\nu)+\iint e^{u(\xi,\theta)-a_{\nu}-b_{\nu}}\left[\left(\frac{d\nu_{\varepsilon}}{d\phi}\right)^{1-\alpha}-\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\right]\>d\phi\>d\mu-1\bigg{|}$
But this is continuous w.r.t. $\varepsilon$. Since $\frac{d\nu_{\varepsilon}}{d\phi}$ is uniformly bounded away from 0 and $\infty$ in a neighbourhood of $\varepsilon=0$, we apply the Dominated Convergence Theorem to get $V(\nu_{\varepsilon})\to V(\nu)$. ∎
Proof of theorem˜12.
@@
Proof of theorem˜14.
If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
@@
is not only a valid probability measure, but also $P|_{\xi}=\nu$, $P|_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
Proof of corollary˜15.1.
If $P,P^{\prime}$ both maximize $U$, then $P|_{\xi},P^{\prime}|_{\xi}$ both maximize $f$. Since $f$ is strictly concave, it follows that the marginals are equal: $P|_{\xi}=P^{\prime}|_{\xi}=\nu$. Because the joint must satisfy the MNL formula, it follows $P$ and $P^{\prime}$ are just versions of one another. ∎
@@
As suggested, the Envelope Theorem provides a nice heuristic interpretation for $a_{\nu}(x)$. It is left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [5].
@@
$g(b,\varepsilon)=\alpha D_{KL}(\nu\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
@@
Sinkhorn’s algorithm
In the discrete case, the problem of finding Schrödinger potentials such that eq.˜5 holds is equivalent to the matrix scaling problem. Formally, we can re-write eq.˜5 in the discrete case as
$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$
If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$, the Schrödinger potentials as diagonal matrices, and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$.
Theorem 16 (Sinkhorn’s algorithm).
Set $a_{t}=0$. For $t\geq 0$, set
$a_{t+1}(\theta)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)$
Then:
  1. 1.
  2. 2.
    $D_{KL}(P\|P_{t})\to 0$
  3. 3.
    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
  4. 4.
    $P_{t}\to P$ in total variation
@@
By assumption, $u$ and $\log\frac{d\nu}{d\phi}$ are totally bounded. This then follows from [6, Theorem 6.15]. ∎

### v2

@@ Conclusion
The exchangeability of the integral and the derivative follows from [folland1999real, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$. The result follows by construction of $Y$. ∎
@@
Proof of theorem˜13.
@@
Proof of theorem˜16.
If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma\ll\phi$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
@@
is not only a valid probability measure, but also $P_{\xi}=\nu$, $P_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
Proof of corollary˜16.1.
If $P,P^{\prime}$ both maximize $U$, then $P_{\xi},P^{\prime}_{X}$ both maximize $f$. Since $f$ is strictly concave, it follows that the marginals are equal: $P_{\xi}=P^{\prime}_{X}=\nu$. Because the joint must satisfy the MNL formula, it follows $P$ and $P^{\prime}$ are just versions of one another. ∎
@@
As suggested, the Envelope Theorem provides a nice heuristic interpretation for the potential $a_{\nu}(x)$. We have left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [milgrom2002envelope].
@@
$g(b,\varepsilon)=\alpha D_{KL}(\nu_{\varepsilon}\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu_{\varepsilon}$
@@
Concavity of $V$
Intuition via duality. Let $g(\nu)$ be
$g(\nu)=-\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
Then, the dual is
$g^{*}(\psi)$ $=\sup_{\nu}\int\psi\>d\nu+\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
$=\sup_{\nu}\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)+\psi(\xi)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
$=\sup_{\nu}\int\psi\>d\nu+\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
where $\overline{\nu}\equiv\overline{\nu}_{\psi},\overline{P}\equiv\overline{P}_{\psi}$ denote the optimal choices of $\nu,P$ given $\psi$. Consider a path $\psi_{\varepsilon}=(1-\varepsilon)\psi+\varepsilon H$. Even though the optimal choice $\overline{\nu}$ depends on $\psi$, by the Envelope Theorem,
$\frac{dg^{*}(\psi_{\varepsilon})}{d\varepsilon}=\int H-\psi\>d\overline{\nu}$
Now consider the biconjugate
$g^{**}(\nu)=\sup_{\psi}\int\psi\>d\nu-g^{*}(\psi)$
For $\psi$ to be chosen optimally, it must be the case that for any direction $H$,
$\int(H-\psi)\>d(\nu-\overline{\nu}_{\psi})=0$
from which it follows that $\psi$ is chosen such that $\nu=\overline{\nu}_{\psi}$. With this being the case,
$g^{**}(\nu)=-\int a_{\nu}\>d\nu-\int b_{\nu}\>d\mu=-V(\nu)$
An Approach Without First-Order Conditions
Consider first the outer integrand of $f$, given $\nu$:
$f(\nu)=\int\Lambda_{\nu}(\theta)\>d\mu(\theta)$
$\Lambda_{\nu}(\theta)=\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}d\phi(\xi)\right)=\log\left(\int\exp\left(\underbrace{u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)}_{h_{\nu,\theta}(\xi)}\right)\ d\nu(\xi)\right)$
The Donsker-Varadhan variational formula states that a log-sup-exp is, in fact, a value function of sorts:
$\Lambda_{\nu}(\theta)=\log\left(\int\exp(h_{\theta}(\xi))\>d\phi(\xi)\right)$ $=\sup_{Q_{\theta}\ll\phi}\bigg\{E_{Q}[h_{\nu,\theta}]-D_{KL}(Q_{\theta}\|\nu)\bigg\}$
$=\sup_{Q_{\theta}\ll\phi}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\bigg\}$
Then,
$f(\nu)=\int\sup_{Q_{\theta}\ll\phi}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\bigg\}\>d\mu(\theta)$
Under suitable regularity conditions, Rockafellar’s interchange theorem allows for the exchange of $\sup$ and $\int$. Assuming such conditions hold,
$f(\nu)=\sup_{Q_{\theta}\ll\phi}\bigg\{\iint u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\>d\mu(\theta)\bigg\}$
Define $P(\xi,\theta)=Q_{\theta}(\xi)\otimes\mu(\theta)$. Then, we have
$f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(\nu\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)\bigg\}$
Lastly, how do we know that $P_{\xi}=\nu$? We don’t: because only at the supremum is this true. But to see that this is true at the supremum, decompose the RHS so that $P_{\xi}$ is plugged in everywhere $\nu$ is:
$f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)+(1-\alpha)\log\left(\frac{d\nu(\xi)}{dP_{\xi}(\xi)}\right)\>dP(\xi,\theta)\bigg\}$
$=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)-(1-\alpha)D_{KL}(P_{\xi}\|\nu)\bigg\}$
And so
$\sup_{\nu\ll\phi}f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)\bigg\}$
$=\sup_{\nu\ll\phi}V(\nu)$
which is precisely the DM’s problem.
As a Selection Device for State-Action Rational Inattention
It is a natural to use the state-characteristic model as a selection device when state-action yields non-unique solutions. Let $V_{\alpha}$ denote the value function of the marginal given $\alpha$, and let $v^{*}_{\alpha}=\arg\max V_{\alpha}$, for $\alpha\in(0,1)$. If $\arg\max V_{0}$ is not a singleton – i.e. multiple solutions in the state-action model – then one can use the state-characteristic model as a selection device by setting $\nu_{0}^{*}=\lim_{\alpha\to 0}\nu^{*}_{\alpha}$. Since the optimal marginal $\nu^{*}_{\alpha}$ is unique for each $\alpha$.
If such an limit exists, it turns out to be the minimum divergence marginal:
$\lim_{\alpha\to 0}\nu^{*}_{\alpha}=\operatorname*{arg\,min}_{\nu\in\arg\max V_{0}}D_{KL}(\nu\|\phi)$
which, ex-post, is both unsurprising (entropic regularization naturally leads to selecting for the maximum entropy choice) and a reasonable choice. Therefore, what we propose is not to find $\nu_{\alpha}^{*}$ for a sequence of $\alpha$’s, which would be cumbersome, but to simply select the minimum divergence marginal.
The trick is to write $V_{\alpha}$ in terms of the dual. Instead of optimizing the dual over $a,b$, we plug the Schrödinger equation into the dual and optimize only over $b$. Moreover, since we get one normalization, we restrict $\int b\>d\mu=0$. Further, note that if $a$ satisfies the Schrödinger equation, then
$\int e^{u-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu-1=0$
so from the dual, we get
$V_{\alpha}(\nu)$ $=\inf_{b\ s.t.\int b\ d\mu=0}\int\log\left(\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\right)+\log\left(\int e^{u-b}d\mu\right)\>d\nu$
$=-\alpha D_{KL}(\nu\|\phi)+\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)d\nu$
The thing to note here is that $b_{\nu}$, which minimizes $\int\log\left(\int e^{u-b}d\mu\right)\>d\nu$ subject to $\int b\>d\mu=0$, does not depend on $\alpha$. So, $\alpha\mapsto V_{\alpha}(\nu)$ is simply an affine function with slope $D_{KL}(\nu\|\phi)$ and $y$-intercept $\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\ d\nu$. Then, define
$\alpha\mapsto\sup_{\nu}V_{\alpha}(\nu)$
to be the upper envelope. The pointwise supremum of a family of affine functions is convex (and trivially, continuous). As $\alpha\to 0$, $D_{KL}(\nu^{*}_{\alpha}\|\phi)$ increases, and therefore, if $\lim_{\alpha\to 0}\nu^{*}_{\alpha}$ exists and is in $\arg\max V_{0}$, it must be the element in the argmax which minimizes $D_{KL}(\nu\|\phi)$. In general, for all $\alpha>0$, $D_{KL}(\nu_{\alpha}^{*}\|\phi)\geq D_{KL}(\nu_{0}\|\phi)$ for all $\nu_{0}\in\arg\max V_{0}$.
Lemma 18.
$\arg\max V_{\alpha}$ is convex for $\alpha\in[0,1)$
@@
Two observations: $x\mapsto x^{1-\alpha}$ and $x\mapsto x^{\alpha}$ are concave, and $x\mapsto 1/x$ is convex, for $x\in(0,1)$. This means that for $\nu,\gamma\in\Delta X$,
$\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
$\leq\int\frac{e^{u(\xi,\theta)}}{\beta\int e^{u(x,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(x)+(1-\beta)\int e^{u(x,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
$\leq\beta\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)+(1-\beta)\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
$\leq\beta\left(\frac{d\nu}{d\phi}\right)^{\alpha}+(1-\beta)\left(\frac{d\gamma}{d\phi}\right)^{\alpha}$
$\leq\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{\alpha}$
Thus,
$\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)-1\leq 0$
However,
$\int\underbrace{\left[\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)-1\right]}_{\leq 0}\ d(\beta\nu+(1-\beta)\gamma)(\xi)=0$
which means
$\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)=1$
almost surely. Therefore,
$P(\cdot|\theta):=\int_{\cdot}\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}d(\beta\nu+(1-\beta)\gamma)(\xi)$
is a valid kernel, insofar as $P_{\xi}=\beta\nu+(1-\beta)\gamma$. Now, suppose $\nu,\gamma\in\arg\max V_{0}$. We get
$V(\beta\nu+(1-\beta)\gamma)$ $\geq U(P)$
$=\int u(\xi,\theta)-\alpha\log\left(\frac{d(\beta\nu+(1-\beta)\gamma)(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi|\theta)}{d(\beta\nu+(1-\beta)\gamma)(\xi)}\right)\>dP$
$=\int\log\left(\int e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)\>d\mu(\theta)$
$\geq\int\beta\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)+(1-\beta)\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)\>d\mu(\theta)$
$=\beta V(\nu)+(1-\beta)V(\gamma)$
One final point to make is that the maximum entropy selector tends to select for non-sparse consideration sets – that is, if there are two consideration sets, their union is also a consideration set, and the maximum entropy selector will prefer to select for the union. Depending on the context, either maximum entropy or consideration set sparsity can be reasonable descriptions of weak human preferences, but they are inherently contradictory, so if part of the impetus for using a rational inattention model is to induce sparse consideration sets, then using a maximum entropy selector would be counterproductive in spirit. And it is no more inherent to use maximum entropy than it is to use minimum cardinality to select among possible outcomes.
Example. Consider a state-action model with a state space
$\Theta=\{1,2,3,4\}$ $\mu(\theta)=1/4$
and an action space
$A=\bigg\{\{1,2\},\{1,3\},\{1,4\},\{2,3\},\{2,4\},\{3,4\}\bigg\}$
and a utility
$u(a,\theta)=\textbf{1}\{\theta\in a\}$
That is, the DM chooses a pair of states, and gets utility $1$ iff the drawn state corresponds to one in the chosen pair.
It is easy to see that any pair of actions whose disjoint union is $\Theta$ is a valid consideration set. One need only check the necessary and sufficient condition provided by [caplin2019rational]. Any convex combination of optimal marginals is also an optimal marginal given that $V$ is concave, meaning a union of two optimal consideration sets is an optimal consideration set. So, $A$ itself is a valid consideration set. What is not a valid (optimal) consideration set, however, is something like $\bigg\{\{1,2\},\{2,3\},\{3,4\}\bigg\}$. [^7] It is easy to show that if $\nu(2,3)>0$, then $\sum_{\theta}\frac{e^{u((2,3),\theta)}}{\sum_{a}\nu(a)e^{u(a,\theta)}}\mu(\theta)>1$ The idea is that if the consideration set was $\bigg\{\{1,2\},\{2,3\},\{3,4\}\bigg\}$ then receiving the action recommendation signal $\{1,2\}$ provides additional but not-utility-relevant information about whether the true state is $1$. In an optimal information policy, the action recommendation $(i,j)\in A$ should not differentiate (reveal unnecessary information) about whether the true state is $i$ or $j$, because once the DM chooses $(i,j)$, such information would be irrelevant. To see this in an extreme case, suppose $\nu(2,3)=0.99$. Then, receiving the action recommendation $(1,2)$ would strongly indicate that the true state is $1$.
The point is that there are sparse optimal consideration sets like $\bigg\{\{1,2\},\{3,4\}\bigg\}$ and non-sparse ones, like $A$. The maximum entropy selector, however, will choose the non-sparse one, assuming that $\phi(a)=1/6$. And indeed, it will be the case $\nu^{*}_{0}(a)=1/6$.
Fixed-Point Iteration
Let $X$ be discrete. Let $\delta_{x}$ denote the Dirac measure at $x\in X$ and let $\partial_{x}f(\nu)$ be the Gateaux derivative in the direction of $\delta_{x}$. Again, let $g=d\nu/d\phi$, generically. We abuse notation by using $g$ and $\nu$ interchangeably: we use $f(g)$ to mean $f(\nu)$, $\partial_{x}f(g)$ to mean $\partial_{x}f(\nu)$, $T\nu$ to mean $Tg$, etc. It should be clear in each instance whether we are referring to the density or the measure.
Lemma 19.
Let $k\geq 1-\alpha$ and $n\geq 0$. Then,
$\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{k}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n}\>d\phi(\xi)\right)\>d\mu(\theta)$
$\leq(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)\>d\mu(\theta)$
with equality only if for $\mu$-a.e. $\theta$,
$e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}\text{ is linearly dependent on }e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\text{ in }L^{1}(\phi)$
Proof.
We exploit the fact that
$\partial_{\xi}f(g)=(1-\alpha)\left(\frac{Tg(x)}{g(x)}-1\right)$
Re-arrange to get
$g(x)=\frac{Tg(x)}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}$
We can re-write $g(\xi)$ as
$g(\xi)^{k}=g(\xi)^{k-(1-\alpha)^{2}}\left(\frac{Tg(\xi)}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{(1-\alpha)^{2}}$
Plugging this in, we get
$\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{k}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n}\>d\phi(\xi)\right)\>d\mu(\theta)$
$=\int\log\left(\int e^{\alpha u(\xi,\theta)}e^{(1-\alpha)u(\xi,\theta)}[Tg](\xi)^{(1-\alpha)^{2}}g(\xi)^{k-(1-\alpha)^{2}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n+(1-\alpha)^{2}}\>d\phi(\xi)\right)\>d\mu(\theta)$
$\leq\int\log\left[\left(\int e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}d\phi(\xi)\right)^{1-\alpha}\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)^{\alpha}\right]\>d\mu(\theta)$
$=(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)\>d\mu(\theta)$
via Holder’s inequality. ∎
Starting with $k=1-\alpha$, we see that $\frac{1-\alpha-(1-\alpha)^{2}}{\alpha}=(1-\alpha)\left(\frac{1-(1-\alpha)}{\alpha}\right)=1-\alpha=k$. Starting from $n=0$, by iteratively adding $(1-\alpha)^{2}$ and dividing by $\alpha$, we get convergence to
$(1-\alpha)^{2}\left(\frac{1}{\alpha}+\frac{1}{\alpha^{2}}+\frac{1}{\alpha^{3}}+\cdots\right)=(1-\alpha)^{2}\left(\frac{1}{1-\alpha}-1\right)$
Thus,
$f(g)$ $=\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{1-\alpha}\>d\phi(\xi)\right)\ d\mu(\theta)$
$\leq(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{1-\alpha}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{(1-\alpha^{2})\left(\frac{1}{\alpha}\right)}\ d\phi(\xi)\right)\ d\mu(\theta)$
$\ \vdots$
$\leq(1-\alpha)\left(1+\alpha+\cdots+\alpha^{k}\right)f(Tg)+\alpha^{k+1}\int\log\left(\int\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\left(1+\frac{\partial_{\xi}f(g)}{1-\alpha}\right)^{(1-\alpha)^{2}\left(\frac{1}{\alpha}+\cdots+\frac{1}{\alpha^{k+1}}\right)}}\ d\phi(\xi)\right)\>d\mu(\theta)$
$\ \vdots$
$\leq f(Tg)$
Note that the inequalities, which come from Holder, are usually strict. In order for the inequality to not be strict, $e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}$ must be $\mu$-almost surely linearly dependent on $\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\left(1+\frac{\partial_{\xi}f(g)}{1-\alpha}\right)^{(1-\alpha)^{2}\left(\frac{1}{\alpha}+\cdots+\frac{1}{\alpha^{k+1}}\right)}}$. Even if this holds for some $k$, it won’t hold for $k+1$ unless $\partial_{\xi}f(g)=0$, either of which suffice for optimality.